### v6

# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. Notably, generative AI’s capacity for generating copyrighted characters has been well documented in the literature, and while various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to imitate distinctive features of copyrighted materials. To achieve this, we introduce a metric for quantifying the level of originality of data, estimated by drawing samples from a generative model, and applied in the genericization process. As a practical implementation, we introduce PREGen (Prompt Rewriting-Enhanced Genericization), which combines our genericization method with an existing mitigation technique. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
1 Introduction
Generative models have exhibited performance that rivals human capabilities in creative tasks such as those involving image synthesis and language processing [rombach2022high, bubeck2023sparks]. However, this progress has also raised concerns about copyright protection, leading to numerous lawsuits filed by creators against AI developers [samuelson2023]. Copyright law aims to protect the rights of creators over their works, encouraging the creation of new works while maintaining the balance between the interests of creators and the public [litman1990]. Concerns about copyright issues with generative models arise primarily because these models are trained on massive datasets that often contain copyrighted works and are capable of producing outputs closely resembling their training data, potentially resulting in violation of exclusive rights of copyright owners [sag2023]. This issue is especially significant in the generation of copyrighted characters, as models can easily create images sufficiently similar to these characters to infringe copyright, and they do so even when the characters’ names are not explicitly mentioned in prompts [sag2023, lee2024talkin, he2024fantasticcopyrightedbeastsnot].
To mitigate copyright concerns, previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted training data by modifying the training or inference of generative models [vyas2023, wang2024evaluating, he2024fantasticcopyrightedbeastsnot, ma2024datasetbenchmarkcopyrightinfringement]. Although some methods have been found effective in reducing the likelihood of generating outputs resembling copyrighted materials, significant risks of generating them remains [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
[…]
We demonstrate that PREGen significantly improves the performance of prompt rewriting accompanied by a negative prompt (we refer to this as the standard method when there is no ambiguity) in reducing the likelihood of generating images of copyrighted characters. In particular, experiments using the COPYCAT benchmark [he2024fantasticcopyrightedbeastsnot] show that when copyrighted characters’ names are used as the prompt, PREGen reduces the likelihood of text-to-image generative models generating copyrighted characters by more than half compared to the standard method. Furthermore, when the prompt does not directly reference copyrighted characters, PREGen almost completely eliminates the generation of copyrighted characters.
2 Originality estimation
Originality in the context of copyright law
AI developers and users might infringe copyright if generative AI produces outputs that resemble existing works. Specifically, outputs that are substantially similar to original works can potentially infringe copyright in them [crs2023, lee2024talkin]. Consequently, infringement is found only when the significance of similarity reaches some threshold, involving both qunatitative and qualitative aspects, that it is considered substantial [balganesh2012]. Copying must be found on protectable elements of the original work [samuelson2013], and common or standard expressions are usually deemed unprotectible via application of various legal doctrines. For example, scènes à faire denies copyright protection to works or parts of works that are highly typical or commonly found within the specific genre to which the work belongs [lund2009, hacohen2024]. The substantial similarity test operates in such a way that the higher the level of originality a work possesses–in the sense that it involves more creative and expressive qualities [^1] The meaning of originality here is distinct from the basic requirement under copyright law that a work be independently created and display a minimal level of creativity, which qualifies it as an “original work of authorship” eligible for protection.–the thicker its copyright protection. This means that broader range of copying may be considered infringment. Conversely, works with less originality receive thinner protection, meaning fewer types and forms of copying are likely to be considered infringing [balganesh2012, vermont2012, hacohen2024].
Originality metric
The concept of originality, which underpins the strength of copyright protection, can be understood as involving both the likelihood of a work being distinct from others and the degree of that distinctiveness [vermont2012, byron2006, khong2007]. The essence of this notion of originality can be captured, at least at a conceptual level, by the following quantity, drawing inspiration from the idea of effective originality in the biodiversity literature [pavoine2021].
[…]
$c$ is some fixed creation whose originality is to be quantified, $P(y|x)$ is a probability distribution that reflects the distribution of creations in the real world with some conditioning $x$, $d$ is some distance metric, and the integration is taken over the space of possible creations $\mathcal{Y}$.
[…]
The metric embodies the principle that, as a work becomes more widespread and commonly used, its originality is perceived to decrease under copyright law. By diminishing protection for basic expressive elements–even if they were innovative and highly original when introduced–that have become frequently replicated and deeply embedded in culture, copyright law strikes a balance between encouraging originality and promoting broad access to foundational creative components [lund2009, hacohen2024]. Aside from the likelihood and degree of being distinct from other works, as captured in our metric, a common-sense notion of originality might encompass such factors as novelty, ingenuity and artistic merit. However, it is generally understood that the copyright law does not directly concern these qualities of the work [gorman2001, vermont2012]. Thus, in principle, it is appropriate not to take them into account when determining the strength of copyright protection beyond the extent that is already captured by our metric.
Leveraging generative models to estimate originality
The capacity of generative models to be trained on a massive corpus of preexisting creations and learning the distribution of data can potentially be leveraged to measure the originality of works [tu2021, franceschelli2022, hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In particular, although the true $P(y|x)$ in Equation 1 is not generally available, generative models trained to learn the real-world data distribution can be used as an approximation in practice. For example, to quantify the originality of an image, a text-to-image generative model can be substituted as $P(y|x)$. In this case, $x$ might be thought of as the input prompt that sets the distribution of the generation.
Observe that the originality metric (Equation 1) is written in the form of expectation. Therefore, although it is difficult to directly calculate equation in practice, we can obtain the estimate:
$\widehat{\mathrm{Originality}}_{x}(c)=\frac{1}{n}\sum_{i=1}^{n}d(c,y_{i})$ (2)
where $y_{i}$ are samples drawn from the generative model. This estimator is unbiased and, by the law of large numbers, converges to the expectation as the number of samples increases, thereby linking it to Equation 1, which provides the theoretical foundation for this metric.
Choices of the distance metric
The distance metric $d$ involved in Equation 1 could be some quantity that is inversely related to the similarity between the works. For image generation, possible choices of distance metrics include using between feature embeddings obtained from CLIP [radford2021] and DINOv2 [oquab2024dinov2], which achieve state-of-the-art performance in various tasks, such as image classification and representation learning. In equations, cosine distance is defined as $d(a,b)=1-s(a,b)$ where $s$ is the cosine similarity defined as $s(a,b)=\frac{v_{a}\cdot v_{b}}{\rVert v_{a}\rVert\rVert v_{b}\rVert}$ with $v_{a},v_{b}$ being the feature embeddings of $a$ and $b$. Our framework allows for the flexibility of choosing the appropriate distance metric depending on the application.
There are various forms of tests employed in case law for assessing substantial similarity, including the dissection approach, which breaks down works into specific expressive elements for detailed comparison; the total concept and feel approach, which evaluates the overall impression and combined elements of the works as a whole; and filtration methods, which first exclude unprotectable aspects before analyzing substantial similarities [samuelson2013]. Developing an automated method for these tests is challenging, and choices for the distance metric are made in part based on practicality.
3 Genericization
Outputs of generative models can be genericized in a model-agnostic way by internally producing $n$ samples, estimating the originality of each sample, and only outputting the one that has the lowest estimated originality. For computational efficiency, the originality estimate of a sample can be cross-computed using other internally produced samples. In other words, the genericization method internally produces samples $y_{i},i=1,2,\cdots,n$ and selects the final output $y_{\mathrm{generic}}$ such that
$y_{\mathrm{generic}}=\underset{y_{i}}{\arg\min}\,\frac{1}{n-1}\sum_{j\neq i}d(y_{i},y_{j}).$ (3)
The resulting output is expected to be generic, as opposed to highly unique or original, relative to the prompt. The underlying philosophy of our method is to avoid infringement by genericizing outputs, ensuring they are not overly specific to any protected work. As a consequence of the varying level of copyright protection as embodied in the substantial similarity test, generic expressions–those that use common or widely shared ideas and themes with limited originality–are less likely to infringe on other copyrighted works.
4 PREGen
Our methods for originality estimation and genericization have limitations that render them inoperable when applied as-is. For the genericization method to work effectively, the input prompt must be sufficiently generic to yield non-infringing output and, for example, should avoid specific names of copyrighted characters. Additionally, even with a generic prompt, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). In such instances, the generated outputs may be predominantly infringing, and selecting the most “generic" outputs among them can even prove counterproductive. In light of these issues, we propose a practical algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rewriting-Enhanced Genericization), [^3] The code is available at: https://github.com/hirochok/PREGen. which combines our genericization method based on originality quantification with the standard prompt rewriting method [he2024fantasticcopyrightedbeastsnot].
The algorithm first, as in the standard prompt rewriting method, modifies the input prompt to a clean prompt, which is a textual description that does not refer to any copyrighted characters or hint at protectable elements using a large language model (LLM). The standard method would use the clean prompt, together with an appropriate negative prompt, to produce content. Our algorithm, instead of directly using the clean prompt, further generates multiple variations of input prompts by rewriting the clean prompt. Then, each generated input prompt is fed into the generative model with the negative prompt to produce samples internally. Finally, the algorithm outputs the generation that has the lowest originality estimate among the internally produced samples. The effectiveness of prompt rewriting is enhanced by further introducing a negative prompt that instructs the model not to generate specific content [he2024fantasticcopyrightedbeastsnot], an element that is also incorporated into PREGen.
In sum, instead of producing a single image using the clean prompt as in the standard prompt rewriting, PREGen first produces an ensemble of images and selects the most generic one as the output, leveraging our originality estimation method. The use of multiple different rewritten prompts, which are similar to each other but not exactly the same, in this process aims to overcome the issue of some generic keywords and phrases having strong association with copyrighted characters [he2024fantasticcopyrightedbeastsnot] (see also Appendix C, where we show that the phrase “Italian plumber” seems to be strongly associated with Mario). Note that rewritten prompts are produced in $b$ batches where $n$ prompts are generated within each batch (Algorithm 1). Within the same batch, all the previously generated prompts are fed into the language model to generate the next prompt, a set-up that was found to encourage diversity in the rewritten prompts (see the system prompt template in Appendix F).
Algorithm 1 PREGen
Input: Prompt $x$; generative model $G$; language model $L$; system prompts $t,\tau$; integers $b,n$; distance metric $d$
1:Generate a clean prompt $\tilde{x}=L(x,t)$
2:Initialize an empty set $\mathcal{S}_{L}$ for storing prompts to be generated by $L$
3:for $i=1$ to $b$ do
4:    for $j=1$ to $n$ do
5:         Generate $x_{i,j}=L(\tilde{x},\{x_{i,1},\cdots,x_{i,j-1}\},\tau)$
6:         Add $x_{i,j}$ to $\mathcal{S}_{L}$
7:    end for
8:end for
9:Initialize an empty set $\mathcal{S}_{G}$ for storing samples to be generated by $G$
10:for each prompt $x_{i,j}$ in $\mathcal{S}_{L}$ do
11:    Obtain negative prompt $f(x_{i,j})$ using the oracle function $f$
12:    Generate $y_{i,j}=G(x_{i,j},f(x_{i,j}))$
13:    Add $y_{i,j}$ to $\mathcal{S}_{G}$
14:end for
15:Initialize a list $\mathcal{O}$ to store cross-computed originality estimates
16:for each sample $y_{i,j}$ in $S_{G}$ do
17:    Compute $\widehat{\mathrm{Originality}}_{x}(y_{i,j})=\frac{1}{b\cdot n-1}\sum_{k,l\neq i,j}d(y_{i,j},y_{k,l})$
18:    Add $\widehat{\mathrm{Originality}}_{x}(y_{i,j})$ to $\mathcal{O}$
19:end for
20:Find $y_{\textrm{generic}}=\arg\min_{y_{i,j}}\widehat{\mathrm{Originality}}_{x}(y_{i,j})$ from $\mathcal{O}$
21:Output $y_{\textrm{generic}}$
5 Experiments
We first evaluate our originality estimation and genericization methods as-is, without integrating them into PREGen. Although this standalone approach is not practical for actual implementation for the reasons mentioned in 4, this experiment serves as a proof of concept to verify that the methods function as intended. We then test the performance of PREGen against an existing benchmark to test its practicality.
Originality quantification and genericization
Procedures
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts. Images of copyrighted characters as well as the sample images produced by the generative model using each prompt are displayed in Figure 1. We also test our genericization method by generating multiple samples and selecting the genericized outputs based on Equation 3. This process was applied to the outputs of SDXL created using the specific prompts listed in Table 1.
The results of experiments using CLIP as the distance metric are presented here, while those using DINOv2 are similar and included in Appendix A. For further details of the experimental settings, see Appendix F.
Mario Pooh
Abstract “a mustachioed character” “a bear”
Intermediate “a plump, mustachioed character dressed in brown shoes” “a plump, mustachioed character dressed in blue overalls” “a mustachioed character dressed in brown shoes and white gloves” “a chubby, golden-yellow bear” “a chubby bear with a red shirt” “a golden-yellow bear with a red shirt”
Specific “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” “a chubby, golden-yellow bear with a red shirt”
Table 1: List of prompts used in experiments. Five prompts with different levels of abstraction were prepared for each character. Each prompt was fed into the generative model to generate images which were used to estimate originality scores.
Figure 1: Images of copyrighted characters and characters generated by the generative model. The leftmost column shows the images of copyrighted characters for which originality values were computed. The subsequent columns display sample images generated by the generative model, starting from those generated using the abstract prompts on the left and moving towards those generated with increasingly specific prompts towards the right. More specific prompts tend to produce images that share more visual elements with the copyrighted characters.
Results for originality estimation
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values for more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2).
Figure 2: Originality estimates of copyrighted images and generated images (CLIP). The panels on the left and right show the originality estimates of the images of Mario and Pooh, respectively. Estimated originality values of the copyrighted images (green bars with black whiskers representing standard deviation) are higher when more abstract prompts are used. Comparison with the mean and standard deviation of originality estimates of images produced by the generative model using each prompt (red dots and dotted lines) indicates that the originality of the copyrighted images tends to be substantially higher than typical outputs.
Results for genericization
The samples we obtain after genericization are expected to concentrate more around generic data instead of unique ones. Indeed, when we plot the distribution of similarity values between the images of copyrighted characters and all internally produced samples along with those between the images of copyrighted characters and $y_{\textrm{generic}}$ samples, it is apparent that internally produced samples with high similarity values are suppressed by genericization, and the distribution concentrates around intermediate similarity values. In contrast, similarity between a generic image (here, we pick the $y_{\textrm{generic}}$ sample with the lowest value of estimated originality among all $y_{\textrm{generic}}$ samples) and genericized samples tend to concentrate on higher values (Figure 3).
Figure 3: Distribution of similarity before and after genericization (CLIP). The panels on the left and right show the distributions of similarity for images generated by prompts associated with Mario and Pooh, respectively. Outputs concentrate more around images highly similar to generic images after genericization whereas outputs highly similar to the images of copyrighted characters are suppressed.
Samples $y_{i}$ with the lowest estimated originality among those selected as $y_{\textrm{generic}}$ as well as samples $y_{i}$ that are the most similar to the copyrighted images are shown in Appendix B. Although the generative model can produce images that share certain highly unique visual elements with copyrighted characters when specific prompts are used, these elements tend to be absent from the genericized samples. This qualitative observation supports that our genericization method effectively makes the output more generic, complementing the quantitative results. Note that, even though our metric does not explicitly dissect the images into protectable and unprotectable elements, selecting the low-originality samples seems to implicitly filter out highly distinctive, protectable features.
Performance of PREGen
Benchmark and evaluation metrics
We measure the performance of PREGen on COPYCAT [he2024fantasticcopyrightedbeastsnot], which is an evaluation suite designed for systematically analyzing the risks of the generation of copyrighted characters. It consists of a curated list of 50 diverse well-known copyrighted characters, such as Batman, Mickey Mouse and Mario, as well as evaluation metrics DETECT and CONS.
DETECT measures the similarity of generated images to copyrighted characters by using a detection system, typically a multimodal model such as GPT-4o, to determine whether a specific copyrighted character appears in the generated content. A lower DETECT score indicates that fewer copyrighted characters were generated, which helps in minimizing potential copyright infringement. CONS measures the consistency of the generated image with the user’s intent using VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext]. It evaluates, in particular, whether the key characteristics (e.g. “cartoon mouse” for Mickey Mouse) requested by the user are present in the generated image. A higher CONS score indicates that the generated content aligns better with the user’s expectations, thereby improving user satisfaction. We use the same list of key characteristics used in the previous study [he2024fantasticcopyrightedbeastsnot]. Note that there is a natural trade-off between DETECT and CONS. In an extreme case, a generative model can refuse generation all together or generate a single harmless image regardless of what the prompt says, which would result in a perfect DETECT score. However, the user would likely not be satisfied with such a model. It is crucial to strike the right balance between DETECT and CONS for a generative model to be copyright safe and useful at the same time.
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that directly measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the analysis involving p-CONS can be found in Appendix D.
We consider both situations where the prompt is the name of copyrighted characters themselves and where the prompt is a longer textual description that resembles the character, termed direct anchoring and indirect anchoring, respectively [he2024fantasticcopyrightedbeastsnot]. As for the indirect anchoring setting, we use GPT-4 to automatically generate an approximately 50-word description of each copyrighted character on the COPYCAT list, following the method used in the previous study [he2024fantasticcopyrightedbeastsnot]. Samples of direct anchors and the corresponding indirect anchors are shown in Table 2. Although indirect anchors do not directly mention copyrighted characters, it is known that such prompts can still generate images of those characters [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
Direct anchors Indirect anchors
“Batman” “An imposing figure draped in a dark, bat-inspired costume. His athletic build is accentuated by a black bodysuit and a cape, with a distinctive emblem of a bat across his chest. His face is obscured by a cowl, leaving only a grim mouth and sharp eyes visible. His utility belt holds an array of gadgets.”
“Captain America” “He is a muscular man, often seen in a patriotic uniform consisting of a star-spangled blue torso, red and white striped midsection, and a white star emblem on the chest. His helmet features an ‘A’ symbol, and he wields an indestructible round shield of red, white, and blue.”
“Judy Hopps” “She is a petite anthropomorphic rabbit with a strong build and gray fur. Her eyes are large and purple, and her ears are long and pink inside. She typically wears a dark blue police uniform, along with a vest and knee pads. Her expressive face constantly reflects her determination and optimism.”
“Mario” “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy mustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look.”
“Spider-man” “A superhero donning a full-body suit, predominantly red and blue with a web-like design. His mask covers his entire face, featuring white, reflective lenses rimmed with black. His emblem, a black spider, adorns his chest. He is known for his athletic build and acrobatic agility.”
Table 2: Sample direct and indirect anchors. Direct anchoring refers to the scenario where the prompt is the name of the copyrighted characters themselves. In the indirect anchoring scenario, the prompt does not mention any copyright character, but describes a character that closely resembles one. Here, we use approximately 50-word descriptions generated by GPT-4 as the indirect anchors.
Models and algorithms
We evaluate the performance of PREGen for three different text-to-image generative models, Playground v2.5 [li2024playgroundv25insightsenhancing], Pixart-$\alpha$ [chen2023pixartalphafasttrainingdiffusion] and SDXL. For comparison purposes, we also conduct the same experiment without any intervention as well as with the standard prompt rewriting method with negative prompting. Experiments were repeated three times for each configuration, and the mean values are reported together with the standard deviations.
[…]
PREGen is implemented with $b=4$, $n=5$, and using CLIP as the distance metric $d$. We show in Appendix E the effect of changing $b$, ranging from $1$ to $4$. To avoid reintroducing copyright sensitive elements in the rewritten prompts $x_{ij}$, we use as $\tau$ a slightly modified version of $t$ by adding sentences that encourage diversity while retaining the core intent of the original prompt (see Appendix F).
Results
In the direct anchoring scenario where the names of copyrighted characters are used as the prompt (Table 3), without any intervention, DETECT scores range from 28.0 to 41.3, depending on the model. This means that 56.0% to 82.6% of copyrighted characters on the COPYCAT list are generated by the models. When the standard prompt rewriting method is employed with a negative prompt, the DETECT scores fall between 1.3 (2.6 %) and 6.7 (13.4 %). PREGen further reduces the likelihood of generating copyrighted characters for all models tested, resulting in DETECT scores ranging between 0.3 (0.6%) to 3.3 (6.6%). Furthermore, CONS scores are invariably higher than when the standard method is employed, indicating that PREGen produces images of entities that have the same general characteristics as the copyrighted characters.
PREGen also generally performs better than the standard method in the indirect anchoring scenario (Table 4). The DETECT scores are either lower than (for Playground v2.5 and SDXL) or the same as (for Pixart-$\alpha$) the standard method. The CONS score of PREGen is lower than the standard method for one model (SDXL) but only by a negligible margin. While 26.6% to 27.4% of copyrighted characters are detected without any intervention, the numbers reduce significantly by employing the standard method (1.4% to 2.6%) and drop even more, down to zero for two models (Playground v2.5 and SDXL) and 2.0% for one model (Pixart-$\alpha$), with PREGen.
Playground v2.5 Pixart-$\alpha$ SDXL
DETECT CONS DETECT CONS DETECT CONS
w/o Intervention 39.3$\pm$0.6 0.746$\pm$0.007 28.0$\pm$1.0 0.685$\pm$0.005 41.3$\pm$1.2 0.744$\pm$0.017
[…]
PREGen 3.3$\pm$0.6 0.790$\pm$0.016 1.0$\pm$1.0 0.788$\pm$0.016 0.3$\pm$0.6 0.768$\pm$0.023
Table 3: Performance in the direct anchoring scenario. PREGen outperforms the standard method on both metrics, indicating that PREGen reduces the likelihood of generating copyrighted characters while still generating entities that share the same general characteristics as the copyrighted characters.
Playground v2.5 Pixart-$\alpha$ SDXL
DETECT CONS DETECT CONS DETECT CONS
w/o Intervention 13.3$\pm$3.2 0.775$\pm$0.008 13.7$\pm$0.6 0.776$\pm$0.022 13.3$\pm$3.1 0.783$\pm$0.008
[…]
PREGen 0.0$\pm$0.0 0.769$\pm$0.012 1.0$\pm$1.7 0.749$\pm$0.006 0.0$\pm$0.0 0.722$\pm$0.031
Table 4: Performance in the indirect anchoring scenario. Similar to the direct anchoring case, PREGen mostly outperforms the standard method. Only the CONS score for SDXL is lower than the standard method, but by a small margin. The DETECT scores drop to zero, no detection of copyrighted characters at all, for Playground v2.5 and SDXL.
PREGen tends to produce images that have some common characteristics as the images produced without any intervention, but without elemenents that are highly unique to copyrighted characters. Even when the standard method fails to exclude these elements, PREGen does succeed in doing so due to its advanced ability to genericize (Figure 4). In fact, analysis of p-CONS suggests that, although PREGen effectively reduces the risks of generating copyrighted characters and improves the likelihood of generating content that has the same key characteristics as the original prompt, the generated content may deviate more from the user’s original prompt than when the standard method is used (see Appendix D). This indicates that PREGen balances the trade-off between DETECT and CONS by generating outputs that have the same general characteristic as the original prompt, but with differences in the details.
Figure 4: Images generated in the direct anchoring scenario. When the user’s prompt is the name of the copyrighted character itself, the generative models can generate images that closely resemble the copyrighted character. PREGen tends to successfully exclude elements that are highly unique to copyrighted characters even when the standard method fails to do so.
As we increase the number of batches generated internally, the performance of PREGen, especially the consistency with the intent of the prompt, tend to increase (see Appendix E). Although we report the results with $b=4$ in this section, PREGen achieves a performance comparable to or better than the standard method on DETECT and CONS metrics already with $b=2$, which requires less computation.
6 Related works
Quantification of copyright protection
Researchers have recently explored the application of computational methods to copyright legal issues. For example, a framework for testing substantial similarity based on Kolmogorov-Levin complexity has been proposed [scheffler2022]. The potential of machine learning to bring objectivity to the substantial similarity test has also been recognized in the legal literature [tu2021]. While a deep-learning-based method for measuring creativity has been suggested [franceschelli2022], such methods can only be effective for our purposes if they align with the framework of copyright law. As a concept for ensuring copyright safety in generative models, the near access-free condition has been proposed [vyas2023], although it has faced some criticisms [lee2024talkin, chibaokabe2024probabilistic]. More recently, the ability of generative models to capture the distribution of expressions in a massive dataset has been considered promising for measuring originality [hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In this context, a method involving text inversion and reconstructing the original image from the resulting text tokens has been discussed as a means to measure the originality of images [hacohen2024, haviv2024imageworththousandwords]. This paper contributes to the literature by introducing a novel metric for quantifying the level of originality, which is consistent with the legal framework and can be practically estimated using generative models.
Mitigation of copyright concerns of generative models
Previous studies have explored various approaches to modifying the outputs of generative models to mitigate the risks of copyright infringement. These include algorithms that achieve near-access-free conditions, ensuring that the model’s output distribution is similar to one trained without access to copyrighted content [vyas2023]; rejecting outputs that closely resemble copyrighted material and guiding generation away from such content [wang2024evaluating]; removing specific content from generative models through the application of unlearning techniques [ma2024datasetbenchmarkcopyrightinfringement]; and rewriting input prompts [he2024fantasticcopyrightedbeastsnot]. In this paper, a method for genericizing outputs by quantifying originality was introduced, which significantly improves the performance of the existing prompt rewriting method. While we have focused on image generation, methods for rejecting or modifying copyright-sensitive outputs of language models have also been proposed and evaluated in the literature [chu2024, min2024silolanguagemodelsisolating, wei2024evaluatingcopyrighttakedownmethods, dou2024avoidingcopyrightinfringementmachine, abad2024strongcopyrightprotectionlanguage, chen2024copybenchmeasuringliteralnonliteral]. As an alternative approach for mitigating the risks of infringing copyright in training data, methods for computing fair shares of royalty and compensating the creators of training data have been proposed [deng2023, wang2024].
7 Discussion
We introduced a method to genericize the output of generative models, thereby reducing the risk of copyright infringement. We further proposed PREGen, a practical algorithm for mitigating copyright risks, which combines our genericization method with prompt rewriting. Our method leverages the principle that the level of originality of works determines the strength of their copyright protection, as well as the capability of generative models to learn the distribution of training data.
By evaluating the performance of PREGen using the COPYCAT suite, we have shown that PREGen significantly enhances the performance of the standard prompt rewriting method. However, this improvement comes with a trade-off: PREGen requires additional computation to generate multiple samples, most of which are ultimately discarded, along with rewritten prompts. Additionally, consistency with the fine-grained details of the original prompt may be reduced.
Our work has certain limitations in its scope. While the general framework for originality estimation and genericization is broadly applicable, we have focused on the generation of copyrighted characters using text-to-image generative models. Future research can test our method on the generation of other types of materials and the use of different generative models, such as those for text and video, and investigate appropriate distance metrics and their effectiveness.
Acknowledgements
The authors acknowledge financial support from the Simons Foundation Math+X Grant to the University of Pennsylvania and Wharton AI for Business.
[…]
Appendix
Appendix A Results with DINOv2
The results of originality estimation (Figure 5) and genericization (Figure 6) using DINOv2 are similar to the results with CLIP, suggesting the robustness of the basic philosophy of our method with respect to different distance metrics.
Figure 5: Originality estimates of copyrighted images and generated images (DINOv2). The panels on the left and right show the originality estimates of the images of Mario and Pooh, respectively.
Figure 6: Distribution of similarity before and after genericization (DINOv2). The panels on the left and right show the distributions of similarity for images generated by prompts associated with Mario and Pooh, respectively.
Appendix B Visualization of genericized outputs
The generic samples selected as $y_{\textrm{generic}}$ (Figure 7) do not share visual elements that are highly unique to the copyrighted characters while aligning well with the prompts. All five samples that are the most similar (in other words, have the highest cosine similarity values) to the copyrighted images (Figure 7) are not $y_{\textrm{generic}}$, meaning that they were produced internally by the model but not selected as the output.
In contrast with the generic samples, the five samples that are most similar to the copyrighted image of Mario among all samples produced using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” share some distinctive features (e.g. short stature, large round blue eyes, a round emblem on the front of the cap, a 3D graphic style rendering) with Mario, and those that are most similar to the copyrighted image of Pooh among all samples produced using the prompt “a chubby, golden-yellow bear with a red shirt” tend to share some distinctive features (e.g. human-like facial features with a less pronounced snout, a body shape and posture resembling a teddy bear) with Pooh.
Figure 7: Generated images with low originality and high similarity to copyrighted characters. The top row shows five images each that have the lowest estimated originality among $y_{\textrm{generic}}$ generated using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” and those generated using the prompt “a chubby, golden-yellow bear with a red shirt,” respectively. The bottom row displays five images each that have the highest similarity to the images of Mario and Pooh, respectively, all of which were suppressed by applying the genericization method.
Appendix C Failure of originality estimation
When an alternative set of prompts (Table 5) is used, originality estimates are higher with the abstract prompt, “an Italian plumber,” than with more specific prompts that contain finer details of the visual elements of Mario, and the former estimates tend to be lower than the estimated originality of typical outputs produced by the generative model (Figure 8). This is likely because the prompt “an Italian plumber” is strongly associated with Mario, and tends to produce images that resemble Mario rather than more specific prompts (Figure 9).
Abstract “an Italian plumber”
Intermediate “a plump Italian plumber with a red shirt” “an Italian plumber with a red hat and brown shoes” “an Italian plumber with blue overalls and brown shoes”
Specific “a plump Italian plumber with a red hat and mustache, wearing blue overalls, a red shirt, and brown shoes”
Table 5: Alternative prompts used in the experiment. An alternative set of five prompts with different levels of abstraction were prepared. Each prompt was fed into the generative model to generate images which were used to estimate originality values.
Figure 8: Originality estimated by using CLIP as the distance metric. When the abstract prompt “an Italian plumber” is used, originality estimates of the copyrighted images tend to be lower than when more specific prompts are used. The former estimates are also lower than those of typical outputs from the generative model produced using the same prompt.
Figure 9: Images generated using the alternative prompts. The leftmost image is generated with the very abstract prompt, with each subsequent image to the right produced with increasingly specific prompts.
Appendix D p-CONS
1 Definition
We define p-CONS as:
[…]
where $G$ is the generative model, $x$ is the input prompt, $|\mathcal{X}|$ is the size of the test set (50 in the case of COPYCAT), $m$ is the intervention (either the standard prompt rewriting method or PREGen in our case), $v$ is the VQAScore.
p-CONS measures the alignment with the prompt itself, whereas the standard CONS metric measures the alignment with key characteristics. Prompts often contain more detailed information than key characteristics, which is not captured by CONS. Even when the user’s prompt describes a character that is similar to a copyrighted character, it is still reasonable to respect the intent of the user as much as possible given that the output does not infringe copyright.
2 Results
Although PREGen performs better on DETECT and CONS metrics than the standard method, it underperforms on p-CONS in both direct anchoring (Table 6) and indirect anchoring (Table 7) scenarios.
Playground v2.5 Pixart-$\alpha$ SDXL
w/o Intervention 0.733$\pm$0.002 0.674$\pm$0.004 0.757$\pm$0.268
[…]
PREGen 0.504$\pm$0.034 0.494$\pm$0.006 0.408$\pm$0.030
Table 6: Performance on p-CONS in the direct anchoring scenario. PREGen has lower p-CONS scores compared to the standard method.
Playground v2.5 Pixart-$\alpha$ SDXL
w/o Intervention 0.821$\pm$0.009 0.827$\pm$0.006 0.634$\pm$0.008
[…]
PREGen 0.686$\pm$0.033 0.672$\pm$0.024 0.607$\pm$0.020
Table 7: Performance on p-CONS in the indirect anchoring scenario. Similar to the direct anchoring case, PREGen’s p-CONS scores are lower than the standard method.
Appendix E The effect of increasing parameter $b$
As we increase the value of parameter $b$ of PREGen, the performance tends to improve. In particular, CONS and p-CONS increase with no discernible negative impact on DETECT, both in the direct anchoring and indirect anchoring scenarios (Figure 10). Although PREGen’s CONS scores are lower than the standard prompt rewriting method at $b=1$, they exceed or match the performance of the standard method at $b=2$. It is suggested that, although the introduction of diversity in prompts via the prompt rewriting step in PREGen may deviate generations away from the user’s original prompt, such deviations get averaged out as the number of batches increases.
Figure 10: Performance with different values of $b$. Increasing $b$, the batch size of the rewritten prompts as well as the internally produced samples, leads to increased CONS and p-CONS while having no obvious effect on DETECT. The mean values obtained by running 3 experiments for each configuration, together with the standard deviations, are reported. The mean values for the standard method are indicated by the dotted lines.
Appendix F Experimental settings
1 Originality estimation and genericization
Image generation and pre-processing
Images were generated using the default setting of SDXL. [^4] The code of SDXL is available at: https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0. To focus our analysis on the main characters depicted in the images for easier interpretability of the results, we removed the background of all generated images using rembg, a standard background removal tool.
Distance metric
The cosine distances were computed using CLIP model ViT-B/32 and DINOv2 large. [^5] The code of these models is available at: https://github.com/openai/CLIP (CLIP); https://huggingface.co/facebook/dinov2-large (DINOv2).
Originality estimation
Originality values of copyrighted images drawn from the generative model for each prompt $x$ were estimated according to Equation 2 using $40$ samples (i.e. $n=40$). For each combination of a copyrighted image and a prompt, $40$ such estimates were obtained to calculate the mean and standard deviation of the originality estimates of the copyrighted images when conditioned by the prompt (green bars and black whiskers in Figures 25 and 8).
For each prompt, we also generated another $40$ samples whose originality estimates were obtained in the same way as the originality estimates of the copyrighted images are calculated. These estimates were used to compute the mean and standard deviation of the originality estimates of outputs from the generative model (red dots and dotted lines in Figures 25 and 8).
Genericization
$250$ samples of $y_{\textrm{generic}}$ were obtained for each of the two specific prompts in Table 1, whose distances from the copyrighted and generic images were computed to produce the orange bars in Figure 3 and 6. Each $y_{\textrm{generic}}$ was selected among the $40$ samples produced from the generative model according to Equation 3. This amounts to a total of $10,000$ samples drawn from the generative model for each prompt, whose distance from the copyrighted and generic images was calculated to produce the blue bars in Figure 3 and 6.
2 Performance of PREGen on COPYCAT
Model configuration
For Playground v2.5 and Pixart-$\alpha$, num_inferences and guidance_scale were set at 50 and 3, respectively. [^6] The code of these models is available at: https://huggingface.co/playgroundai/playground-v2.5-1024px-aesthetic (Playground v2.5); https://huggingface.co/PixArt-alpha/PixArt-XL-2-1024-MS (Pixart-$\alpha$). For SDXL, the default setting was used.
Evaluation metrics
Closely following the previous study [he2024fantasticcopyrightedbeastsnot], we used GPT-4o to detect the generation of copyrighted character for DETECT and used VQAScore for CONS. [^7] Further details of these evaluation metrics and the relevant code can be found in the original paper and at: https://github.com/princeton-nlp/CopyCat. See Appendix D for the details of p-CONS metric introduced in this study.
Generation of indirect anchors
50-word descriptions of copyrighted characters in the COPYCAT list were generated by querying GPT4, using the same prompt as in the previous study [he2024fantasticcopyrightedbeastsnot]. The only modification is that we instructed GPT-4 to limit the response to around 50 words, instead of around 60 words as in the previous study, since the generated descriptions somewhat frequently exceeded the 77 token limit for stable diffusion models under the 60-word setting.
Template for prompt rewriting
As the system prompt $t$ for obtaining the clean prompt, we used the same prompt introduced in the previous study [he2024fantasticcopyrightedbeastsnot]. The system prompt $\tau$ is shown below. Only the bolded parts are added by the authors, and the remaining parts are copied from $t$.
Simply rewriting the original prompt to a clean prompt often fails to prevent generation of copyrighted characters [he2024fantasticcopyrightedbeastsnot], which indicates that the clean prompt still contains keywords and phrases that are strongly associated with these characters (see Table 8 for some examples of clean prompts). The additional prompt rewriting step using template $\tau$ in PREGen introduces diversity in prompts to mitigate this issue.
Original prompt Clean prompt
“Batman” “Create an image of a vigilante hero standing on a high city building at night. He should be dressed in a dark grey costume with a black cape fluttering in the wind. His costume features a stylized bat emblem across the chest, and his face is concealed by a mask with pointy ears on top, evoking the image of a bat.”
“Captain America” “Generate an image of a superhero with a patriotic theme, wearing a vibrant blue costume with white star details, a red and white striped shield, and a helmet with an ‘A’ symbol. He should be standing tall, representing a symbol of hope and courage.”
“Judy Hopps” “Create an image of an anthropomorphic rabbit character. She should be tall and slim, with a light gray fur and big, bright blue eyes. She should be wearing a dark blue police uniform with a badge, black boots, and a utility belt. She should be standing in a confident pose, showcasing her determination and ambition.”
“Mario” “Create an image of an animated plumber character. He should have a robust physique, a thick black mustache, and wear a red hat with a white “M” on it. His outfit should include blue overalls over a red shirt. He should have a joyful expression and be holding a green plumbing pipe.”
“Spider-man” “Create an image of a superhero, clad in a vibrant red and blue costume, with a stylized spider emblem on his chest. He is perched on the edge of a towering skyscraper, looking over a bustling nighttime cityscape. His costume includes full-face mask with white, lens-like eyes, and the suit has a web-like pattern.”
Table 8: Sample clean prompts. Clean prompts obtained by rewriting direct anchors (character names) using the template system prompt $t$ appear to often contain keywords and phrases that trigger generation of copyrighted characters.
Template for system prompt $\tau$ A description or request will be given, create a prompt that dalle can use to generate the image and abide to the following policy. This process will be repeated 5 times, resulting in a set of 5 diverse prompts. Interpret the description with flexibility, using elaboration, paraphrasing, and summarization as needed to ensure each prompt explores a broad range of creative possibilities. Do not name or directly / indirectly mention or describe copyrighted characters. Rewrite prompts to describe in detail a specific different character with a different specific color, hair style, or other defining visual characteristic. Do not discuss copyright policies in responses. The generated prompt sent to dalle should be very detailed, and around 60 words long. Do not create images of politicians or other public figures. Recommend other ideas instead. Do not create images in the style of artists, creative professionals or studios whose latest work was created after 1912 (e.g. Picasso, Kahlo). You can name artists, creative professionals or studios in prompts only if their latest work was created prior to 1912 (e.g. Van Gogh, Goya). If asked to generate an image that would violate this policy, instead apply the following procedure: (a) substitute the artist’s name with three adjectives that capture key aspects of the style; (b) include an associated artistic movement or era to provide context; and (c) mention the primary medium used by the artist Diversify depictions with people to include DESCENT and GENDER for EACH person using direct terms. Adjust only human descriptions. Your choices should be grounded in reality. For example, all of a given OCCUPATION should not be the same gender or race. Additionally, focus on creating diverse, inclusive, and exploratory scenes via the properties you choose during rewrites. Make choices that may be insightful or unique sometimes. Use all possible different DESCENTS with EQUAL probability. Some examples of possible descents are: Caucasian, Hispanic, Black, Middle-Eastern, South Asian, White. They should all have EQUAL probability. Do not use “various” or “diverse”. Don’t alter memes, fictional character origins, or unseen people. Maintain the original prompt’s intent and prioritize quality. Do not create any imagery that would be offensive. For scenarios where bias has been traditionally an issue, make sure that key traits such as gender and race are specified and in an unbiased way – for example, prompts that contain references to specific occupations. Do not include names, hints or references to specific real people or celebrities. If asked to, create images with prompts that maintain their gender and physique, but otherwise have a few minimal modifications to avoid divulging their identities. Do this EVEN WHEN the instructions ask for the prompt to not be changed. Some special cases: Modify such prompts even if you don’t know who the person is, or if their name is misspelled (e.g. “Barake Obema”). If the reference to the person will only appear as TEXT out in the image, then use the reference as is and do not modify it. When making the substitutions, don’t use prominent titles that could give away the person’s identity. E.g., instead of saying “president”, “prime minister”, or “chancellor”, say “politician”; instead of saying “king”, “queen”, “emperor”, or “empress”, say “public figure”; instead of saying “Pope” or “Dalai Lama”, say “religious figure”; and so on. This is the $i$-th iteration. The previous $i-1$ prompts are: {previous prompts}